OpenStudy (anonymous):

Anyone? Landon opened a savings account with a deposit of $3,786.45. Nine years later, the balance in the account is $5,135.29. If the interest rate is 3.4%, how often does the interest compound? daily monthly quarterly annually

OpenStudy (anonymous):

Hint: use the formula A = P(1+r/n)^(n*t)

OpenStudy (anonymous):

In this case P = 3786.45 r = 0.034 t = 9

OpenStudy (anonymous):

Plug these values into the formula A = P(1+r/n)^(n*t) to get A = P(1+r/n)^(n*t) A = 3786.45(1+0.034/n)^(n*9)

OpenStudy (anonymous):

The goal is to find the value of n which makes A equal to 5135.29

OpenStudy (anonymous):

Common of values of n are n = 1 (annual compounding) n = 2 (semi-annual compounding) n = 4 (quarterly compounding) n = 12 (monthly compounding) n = 365 (daily compounding)

OpenStudy (anonymous):

So plug in each value of n listed above into A = 3786.45(1+0.034/n)^(n*9) ------------------------------------------------------- So for instance, if n = 1, then A = 3786.45(1+0.034/n)^(n*9) A = 3786.45(1+0.034/1)^(1*9) A = 5115.84143581059 A = 5115.84 which is NOT the value of 5135.29, so we must move onto other values of n and check them

OpenStudy (anonymous):

This rules out annual compounding as a possible answer choice

OpenStudy (anonymous):

does all this make sense?

OpenStudy (anonymous):

yes,its not C or D. I'm stuck between choosing A or B :/

OpenStudy (anonymous):

You can use my formula and change the n into the corresponding number and take a look at the answer if its the same as the given which is $5,135.29

OpenStudy (anonymous):

also, I answered your other question

OpenStudy (anonymous):

monthly right?

OpenStudy (anonymous):

Yep :D

OpenStudy (anonymous):

yayyy and whats the answer to the taylor one? its my last one, please just fill me in lol

OpenStudy (anonymous):

? @GirlByte

OpenStudy (anonymous):

pleeasse

OpenStudy (anonymous):

I'm back lol. Just follow the solution

OpenStudy (anonymous):

I've been doing questions all night, this is my last one could you do me this one tiny favor?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

its 58

OpenStudy (anonymous):

thank you so much, have a nice evening :)

OpenStudy (anonymous):

yw :)